3. A force (32i +173)N acts on a particle. The particle is displaced from 7-(42i+3.7j)m to 11.7i -2.1j)m. Please find the following: A. The displacement, AF, of the particle. B. The work, W, done by the force, F, on the particle.
Answer #1

(A) The displacement vector is

r = rf-ri = (11.7 _ 4.2)i + (-2.1-3.7)J 7.5i-5.8j

(B) The work done is

W = F.r = (32\hat{i} + 17\hat{j}).(7.5\hat{i}-5.8\hat{j}) = 240-98.6=141.4J